# Effect of Synthetic Jets Actuator Parameters on Deep Reinforcement Learning-Based Flow Control Performance in a Square Cylinder

### v2

@@ III.3 Sensitivity analyses of synthetic jets
In the previous section, we observed the suppression process of vortex shedding in the wake of a square cylinder during DRL training at $Re=100$ and $Re=500$. The training results indicate that when the synthetic jets are positioned near the trailing edge corners of the square cylinder, DRL-based AFC can completely suppress vortex shedding in the wake at both $Re=100$ and $Re=500$. Additionally, Chen et al.Chen et al. (2023) and Yan et al.Yan et al. (2023) analyzed the impact of synthetic jet positioning on flow control efficacy, yielding some intriguing conclusions. Building on these findings, we integrate the optimization of jet actuator positioning with the optimization of feedback laws using DRL. This study tests Reynolds numbers of 100 and 500, representing two distinctly different vortex shedding modes. Firstly, we perform flow control using symmetric zero-mass-flux jets positioned at the leading and trailing edges of the square cylinder. The objective is to investigate how the positioning of jets actuators influences flow control performance. Secondly, we design synthetic jets with four different widths to study how the width of the synthetic jets affects flow control performance.
@@
Chen et al.Chen et al. (2023) 100 Leading NO 1.480 1.370 7.20 - - - 0.530 0.190 64.2
Chen et al.Chen et al. (2023) 100 Trailing YES 1.480 1.280 13.3 - - - 0.530 0.080 84.9
Our study 100 Leading NO 1.549 1.353 12.7 2.020 0.001 99.9 0.179 0.033 81.5
Our study 100 Trailing YES 1.549 1.325 14.4 2.020 0.001 99.9 0.179 0.025 86.1
Yan et al.Yan et al. (2023) 500 Leading NO 1.980 1.100 44.4 0.470 0.070 84.3 1.260 0.170 85.9
Yan et al.Yan et al. (2023) 500 Trailing NO 1.980 1.950 1.40 0.470 0.520 10.7 1.260 1.070 13.2
Our study 500 Leading NO 2.060 0.711 65.5 0.440 0.268 86.7 1.173 0.650 44.4
Our study 500 Trailing YES 2.060 1.001 51.4 0.441 0.042 90.5 1.173 0.294 74.9
As illustrated in Table. 3, for $Re=100$, positioning the synthetic jets near the leading edge of the square cylinder led to a 12.7% reduction in the average $C_{D}$ compared to the baseline. The standard deviation of $C_{D}$ was suppressed by 99.9%, and the standard deviation of the $C_{L}$ was controlled by 81.5%. However, the wake remained unstable. This instability is evident in the oscillations observed in the $C_{D}$ and $C_{L}$ time history curves of the controlled flow, as depicted in Fig. 13. When the jets position are designed at the front corner points, the controlled flow exhibits significant reductions in both $C_{D}$ and $C_{L}$ compared to the baseline. However, when compared to the jet position designed at the trailing corner points, fluctuations in the $C_{D}$ and $C_{L}$ of the controlled flow still persist. When positioning the synthetic jets near the trailing edge of the square cylinder, a 14.4% decrease in the average $C_{D}$ was observed compared to the baseline. The standard deviation of $C_{D}$ was suppressed by 99.9%, and the standard deviation of the $C_{L}$ was controlled by 93.1%. In addition, when the synthetic jets are positioned at the trailing edge corners of the square cylinder, the vortices in the controlled flow are completely suppressed, as detailed in Fig. 11. As shown in Fig. 13, the controlled flow exhibited a significant reduction in both $C_{D}$ and $C_{L}$, with no fluctuations. Therefore, for $Re=100$, positioning the synthetic jets near the trailing edge is advantageous for reducing the average $C_{D}$ and fully suppressing vortex shedding. Chen et al.Chen et al. (2023) obtained similar results at $Re=100$, indicating that vortex shedding is completely suppressed only when the synthetic jets are positioned near the trailing edge of the square cylinder. Additionally, the reduction in $C_{D}$ is more pronounced when synthetic jets are placed near the trailing edge.
@@
Figure 13: Effect of synthetic jets positioning on flow control at $Re=100$ and $Re=500$. The leading jets are positioned near the front corners of the square cylinder, while the trailing jets are positioned near the trailing corner points. The precise locations of the synthetic jets are detailed in Fig. 2. (a) $Re=100$, $C_{D}$; (b) $Re=100$, $C_{L}$; (c) $Re=500$, $C_{D}$; (d) $Re=500$, $C_{L}$.
In the case of $Re=500$ in Table. 3, when the synthetic jets are positioned near the front corner, the $C_{D}$ reduction rate reaches 65.5%, with a decrease of 86.7% in the std of the $C_{D}$ and a 44.4% reduction in the std of the $C_{L}$. When the synthetic jets are placed near the trailing corner, the $C_{D}$ reduction rate is 51.4%, with a decrease of 90.5% in the std of the $C_{D}$ and a 74.9% reduction in the std of the $C_{L}$. Therefore, when $Re=500$, setting the position of the synthetic jet near the front corner point achieves a higher drag reduction rate than that near the trailing corner point. In the study by Yan et al.Yan et al. (2023), it was similarly found that positioning the jets near the front corner points of the square cylinder is more advantageous for reducing the drag coefficient. Although positioning the synthetic jets near the leading corner results in a higher reduction rate of $C_{D}$, the flow stability of the controlled flow is stronger when the synthetic jets are positioned near the trailing corner. As shown in Fig. 13, both $C_{D}$ and $C_{L}$ in the controlled flow reach a stable state when the synthetic jets are placed near the trailing corner. On the contrary, when the synthetic jets are positioned near the leading corner, both $C_{D}$ and $C_{L}$ exhibit periodic oscillations.
@@
As an illustration at $Re=500$, we present the streamline visualizations for the baseline flow, controlled flow with the synthetic jets positioned near the leading edge, and controlled flow with the synthetic jets positioned near the trailing edge. Both the baseline and controlled flows represent the stable flow states attained after sufficient flow development. In Fig. 14, streamline plots are provided for these three scenarios, along with magnified views focusing on the flow patterns around the square cylinder. As shown in Fig. 14(a), flow separation occurs at the leading edge of the square cylinder in the baseline flow, forming recirculation regions on both sides of the cylinder, with significant oscillations in the wake flow streamlines. Fig. 14(b) illustrates the streamlines around the square cylinder, along with separation bubbles near the sidewalls of the cylinder. Fig. 14(c) illustrates the stabilized flow field achieved after implementing control with the synthetic jets positioned near the leading edge of the square cylinder. In Fig. 14(d), the placement of the synthetic jets and the resulting streamlines around the cylinder are shown. On one side of the cylinder, the influence of the jet’s blowing action enlarges the flow separation region, even forming a smaller separation bubble ahead of a larger one. Conversely, on the other side of the cylinder, the suction action of the jet creates a relatively smaller separation bubble near the leading corner. The separated flow near the leading edge quickly reattaches to the sidewall, undergoes flow separation again at the trailing edge, and sheds a vortex. Under the control of the synthetic jets, the near-field vortices continuously shed alternately, while the far-field wake gradually stabilizes.
@@
Moreover, the synthetic jets are positioned near the trailing corner of the square cylinder, resulting in a stabilized wake flow after control in Fig. 14(e). Fig. 14(f) illustrates the placement of the synthetic jets and the streamlines around the square cylinder. Due to the influence of the synthetic jets, a stable and symmetric separation bubble forms on the leeward side of the cylinder, with minimal oscillation observed in the wake region’s streamlines. Under the flow conditions at $Re=500$, the stability of the controlled flow is effectively enhanced when the synthetic jets are positioned near the leading edge. When the synthetic jets are placed near the trailing edge, the controlled flow achieves a stable state without vortex shedding. In the work of Yan et al.Yan et al. (2023), the DRL control framework they designed focuses on reducing $C_{D}$ and $C_{L}$, and the control strategy produces obvious control effects at multiple Reynolds numbers. In our research, we not only focus on reducing the lift and drag coefficients, but also on controlling the stability of the wake region and the detachment of the vortex behind the square cylinder. The control strategy we obtained completely suppresses the shedding of vortices in the square cylinder wake area under the flow of $Re=500$. This discrepancy arises because, in our research, the agent is capable of observing not only the flow field information around the square cylinder but also, more importantly, the flow field information in the most pulsating regions of the wake field. Li and ZhangLi and Zhang (2022) employed a similar probe design approach and developed a control strategy capable of completely suppressing the Kármán vortex street phenomenon in the wake field of a cylinder.
To comprehensively illustrate the control process of synthetic jets positioned near the front and trailing corner points on the wake flow field around a square cylinder, we present the case of $Re=500$. We display the contours of instantaneous vorticity from the initiation to the ending of the control process. Initially, the activation of the synthetic jets affects only the flow near the square cylinder. We focus on this region initially, and as the control progresses, the influence extends to the wake field, ultimately showcasing the instantaneous vorticity values across the entire computational domain. When the synthetic jets are positioned near the leading edge of the square cylinder, the evolution of the instantaneous vorticity field during the control process is described using five representative time snapshots, as depicted in Fig. 15. In Fig. 15(a), the control action initiates near the leading edge of the square cylinder, causing a slight disruption in the originally continuous vorticity field. Moving to Fig. 15(b), as the synthetic jets near the leading edge continue their control action, their influence on the vorticity around the square cylinder gradually extends rearward. By the time Fig. 15(c) is reached, vortices generated from flow separation near the leading edge reattach to the cylinder’s sidewalls, elongating along the walls towards the rear corner of the cylinder. As the flow control progresses to Fig. 15(d), the originally continuous large vortex around the cylinder was disrupted and divided into several smaller vortices. These smaller vortices developed towards the rear corner, and after separating again, formed alternating vortex shedding in the wake region. When the synthetic jets were located near the leading edge of the square cylinder, after the flow control reached a stable state in Fig. 15(e), alternating vortex shedding occurred on the leeward side of the cylinder, and the shed vortices gradually dissipated in the far field due to the dissipation effect.
@@
Next, we will investigate the impact of synthetic jets width on flow control performance. To achieve the complete suppression of vortices in the wake flow, we position the synthetic jets near the trailing corner of the square cylinder. The width of the synthetic jets under study ranges from $D/20$, $D/25$, $D/30$ to $D/35$, where D represents the characteristic dimension of the square cylinder. Performance metrics for flow control include the reduction rate of the average drag coefficient, and the standard deviations of lift and drag coefficients, as well as the average and standard deviation of action. The stability of training results is evaluated using the standard deviations of $C_{D}$, $C_{L}$, and action. We conduct experiments under two flow conditions, $Re=100$ and $Re=500$, to test the training results for the four jet widths. The summarized performance metrics are presented in Table. 4, and the time-history curves of $C_{D}$, $C_{L}$, and action are illustrated in Fig. 17.
@@
Figure 17: Effect of synthetic jets width on flow control, for $Re=100$ and $Re=500$. $W_{1}$, $W_{2}$, $W_{3}$, and $W_{4}$ represent jet widths of $D/20$, $D/25$, $D/30$, and $D/35$, respectively. (a) $Re=100$, $C_{D}$; (b) $Re=500$, $C_{D}$; (c) $Re=100$, $C_{L}$; (d) $Re=500$, $C_{L}$; (e) $Re=100$, Action; (f) $Re=500$, Action.
@@
Table. 4 presents the results of training the jets at four different widths. At $Re=100$, the average drag coefficient reduction ranges from 13.1% to 15.2% as the width of the synthetic jets transitions from $D/20$, $D/25$, $D/30$ to $D/35$, with the standard deviations of the $C_{D}$ all approaching 99.9%. The impact on the drag reduction rate and the stability of the $C_{D}$ is minor across the four widths of the synthetic jets. However, the reduction in the standard deviation of the $C_{L}$ under $D/30$ and $D/35$ widths is smaller than that under $D/20$ and $D/25$. Further reduction in the width of the synthetic jets adversely affects the stability of the $C_{L}$. The average and standard deviation of actions are greater under $D/30$ and $D/35$ compared to $D/20$ and $D/25$. Decreasing the width of the synthetic jets leads to an increase in both the average and standard deviation of actions utilized. This shows that as the jet width decreases, the control system requires larger amplitude and varying range of actions to achieve the same control effect. At $Re=500$, the reduction rate in the average drag coefficient ranges from 48.9% to 53.0%, with the width of the synthetic jets having minimal impact on the drag reduction rate. However, the reduction rates in the standard deviations of the $C_{D}$ and $C_{L}$ decrease with decreasing width of the synthetic jets. Additionally, the average value and standard deviation of action show consistency with $Re=100$, where narrowing the width of the synthetic jets results in an increase in both the average value and standard deviation of action.
@@
The time curves of $C_{D}$, $C_{L}$, and control actions for the square cylinder at $Re=100$ and $Re=500$ are depicted in Fig. 17. For the $Re=100$ case, at four different synthetic jets widths, the initial stages show significant fluctuations in $C_{D}$, $C_{L}$, and control actions. After this period of intense fluctuation, $C_{L}$ and control actions stabilize near zero, while $C_{D}$ stabilizes near its minimum value. Upon observing the stabilized results, it is evident that the width of the synthetic jets has a minimal impact on drag reduction, although narrower jet widths adversely affect the stability of $C_{L}$ and control actions. In the case of $Re=500$, $C_{D}$, $C_{L}$, and control actions also experience significant fluctuations during the initial stages across all four synthetic jets widths. However, the stability of $C_{D}$, $C_{L}$, and control actions is generally worse compared to the $Re=100$ case. This is attributed to the inherently more unstable flow field at $Re=500$, making control more challenging. Comparing the control results for different synthetic jets widths, further narrowing the jet width has a more significant negative impact on the stability of $C_{D}$, $C_{L}$, and control actions.

### v3

@@ III.3 Parameterization and Sensitivity Analysis of Synthetic Jets
In the previous section, we observed the suppression process of vortex shedding in the wake of a square cylinder during DRL training at $Re=100$ and $Re=500$. The training results indicate that when the synthetic jets are positioned near the trailing edge corners of the square cylinder, DRL-based AFC can completely suppress vortex shedding in the wake at both $Re=100$ and $Re=500$. Additionally, Chen et al.Chen et al. (2023) and Yan et al.Yan et al. (2023) analyzed the impact of synthetic jets positioning on flow control efficacy, yielding some intriguing conclusions. Building on these findings, we integrate the optimization of jet actuator positioning with the optimization of feedback laws using DRL. This study tests $Re$ of 100 and 500, representing two distinctly different vortex shedding modes. First, we perform flow control using symmetric zero-mass-flux jets positioned at the leading and trailing edges of the square cylinder. The objective is to investigate how the positioning of jets actuators influences flow control performance. econdly, we design synthetic jets with four different widths to study how the width of the synthetic jets affects flow control performance. Then, we analyze the impact of jet parameters on control performance by examining the physical mechanisms of flow control achieved through synthetic jets.
III.3.1 Location of Synthetic Jets
In Tab. 3, the training results indicate that at $Re=100$, placing the synthetic jets near the leading edge of the square cylinder reduces the average drag coefficient $C_{D}$ by 12.7% compared to the baseline. The standard deviation of $C_{D}$ is suppressed by 99.9%, and the standard deviation of $C_{L}$ is controlled by 81.5%. When the synthetic jets is positioned near the trailing edge of the square cylinder, the average $C_{D}$ decreases by 14.4% compared to the baseline. The standard deviation of $C_{D}$ is suppressed by 99.9%, and $C_{L}$ is controlled by 93.1%. At $Re=500$, placing the synthetic jets near the leading corner of the cylinder achieves a drag reduction rate of 65.5%. The standard deviation of $C_{D}$ decreases by 86.7%, and $C_{L}$ decreases by 44.4%. Placing the synthetic jets near the trailing corner of the cylinder achieves a drag reduction rate of 51.4% at $Re=500$. The standard deviation of $C_{D}$ decreases by 90.5%, and $C_{L}$ decreases by 74.9%. Our training results demonstrate that at $Re=100$, positioning the synthetic jets at the trailing corner of the square cylinder is more beneficial for reducing drag. However, for $Re=500$, placing the synthetic jets near the leading corner of the cylinder shows a more pronounced drag reduction effect. Regarding the influence of synthetic jets position on drag reduction, the findings of Chen et al. and Yan et al. align with our results.
@@ III.3.1 Location of Synthetic Jets
Chen et al.Chen et al. (2023) 100 Leading NO 1.480 1.370 7.20 % - - - 0.530 0.190 64.2 %
Chen et al.Chen et al. (2023) 100 Trailing YES 1.480 1.280 13.3 % - - - 0.530 0.080 84.9 %
Our study 100 Leading NO 1.549 1.353 12.7 % 2.020 0.001 99.9 % 0.179 0.033 81.5 %
Our study 100 Trailing YES 1.549 1.325 14.4 % 2.020 0.001 99.9 % 0.179 0.025 86.1 %
Yan et al.Yan et al. (2023) 500 Leading NO 1.980 1.100 44.4 % 0.470 0.070 84.3 % 1.260 0.170 85.9 %
Yan et al.Yan et al. (2023) 500 Trailing NO 1.980 1.950 1.40 % 0.470 0.520 10.7 % 1.260 1.070 13.2 %
Our study 500 Leading NO 2.060 0.711 65.5 % 0.441 0.268 86.7 % 1.173 0.650 44.4 %
Our study 500 Trailing YES 2.060 1.001 51.4 % 0.441 0.042 90.5 % 1.173 0.294 74.9 %
The time history curves of the controlled flow’s $C_{D}$ and $C_{L}$ are illustrated in Fig. 12. At $Re=100$, we observe periodic fluctuations in the baseline flow’s lift coefficient. When the synthetic jets are positioned near the leading and trailing corners of the cylinder, both drag and lift coefficients exhibit transient changes upon activation of the synthetic jets, quickly reducing to minimum levels. Notably, positioning the synthetic jets near the trailing corner results in more stable $C_{D}$ and $C_{L}$ coefficients. This difference becomes more pronounced at $Re=500$, where positioning the synthetic jet near the leading corner still results in periodic fluctuations in both coefficients. Conversely, positioning the synthetic jets near the trailing corner stabilizes both $C_{D}$ and $C_{L}$ coefficients. In both $Re=100$ and $Re=500$ cases, positioning the synthetic jets near the trailing corner stabilizes the flow’s $C_{D}$ and $C_{L}$ coefficients effectively. This observation is consistent with the findings in Tab. 3, which describe how only positioning the synthetic jets near the trailing corner suppresses vortex shedding in the wake completely. Similar control effects were observed by Chen et al. at $Re=100$. To investigate the reasons behind these results, we compare streamline plots of the baseline and controlled flows in Fig. 12, providing a detailed analysis of how the position of the synthetic jets influences control effectiveness.
@@
Figure 12: Effect of synthetic jets positioning on flow control at $Re=100$ and $Re=500$. The leading jets are positioned near the leading corners of the square cylinder, while the trailing jets are positioned near the trailing corner points. The precise locations of the synthetic jets are detailed in Fig. 2. (a) $Re=100$, $C_{D}$; (b) $Re=100$, $C_{L}$; (c) $Re=500$, $C_{D}$; (d) $Re=500$, $C_{L}$.
Figure 13: $Re=100$, Streamlines of a square cylinder during baseline and flow control. (a) baseline flow. (b) the controlled flow with the synthetic jets positioned near the trailing edge. (c) for the controlled flow with the synthetic jets positioned near the leading edge.
Fig. 13 illustrates the streamlines of the baseline flow at $Re=100$, as well as the streamlines with synthetic jets positioned near either the leading or trailing corners. In the baseline flow, separation primarily occurs near the trailing corners. When synthetic jets are placed near the trailing corners, performing suction or blowing control directly disrupts the vortices near the trailing corners, delaying the separation point. In the baseline flow, there is no reverse flow on the sidewalls of the square cylinder, and the fluid separates near the trailing corners. However, when synthetic jets are placed near the leading corners, performing suction or blowing control creates recirculation regions on both sides of the cylinder, leading to separation again near the trailing corners. In this scenario, regardless of how the jet speed is adjusted, flow separation reoccurs at the trailing corners.
Our research indicates that for the baseline flow at $Re=100$, positioning the synthetic jets near the trailing corner yields a maximum drag reduction of 14.4%. For the baseline flow at $Re=500$, the optimal drag reduction efficiency is achieved by placing the synthetic jet near the leading corner, with a maximum drag reduction rate of 65.5%. This phenomenon is related to the differing locations of flow separation around the square cylinder at $Re=100$ and $Re=500$. At $Re=100$, flow separation occurs near the trailing corner, whereas at $Re=500$, flow separation occurs near the leading corner, with the separated bubble reattaching to the sidewall and subsequently reattaching near the trailing corner. Based on the flow control effects given by DRL, it is found that placing the synthetic jets at the location of flow separation is more conducive to achieving control objectives. Similarly, KIM2009172 and KIM2009172 observed that when controlling flow separation on an airfoil using synthetic jets, the control effectiveness is optimal when the separation point coincides with the location of the synthetic jets.
@@
As an illustration at $Re=500$, we present the streamline visualizations for the baseline flow, controlled flow with the synthetic jets positioned near the leading edge, and controlled flow with the synthetic jets positioned near the trailing edge. Both the baseline and controlled flows represent the stable flow states attained after sufficient flow development. In Fig. 14, streamline plots are provided for these three scenarios, along with magnified views focusing on the flow patterns around the square cylinder. In Fig. 14(a), flow separation occurs at the leading edge of the square cylinder in the baseline flow, forming recirculation regions on both sides of the cylinder, with significant oscillations in the wake flow streamlines. The streamlines around the square cylinder, along with separation bubbles near the sidewalls, are depicted in Fig. 14(b). The stabilized flow field achieved after implementing control with the synthetic jets positioned near the leading edge of the square cylinder is illustrated in Fig. 14(c). In Fig. 14(d), the placement of the synthetic jets and the resulting streamlines around the cylinder are displayed. On one side of the cylinder, the influence of the jet’s blowing action enlarges the flow separation region, even forming a smaller separation bubble ahead of a larger one. Conversely, on the other side of the cylinder, the suction action of the jet creates a relatively smaller separation bubble near the leading corner. The separated flow near the leading edge quickly reattaches to the sidewall, undergoes flow separation again at the trailing edge, and sheds a vortex. Under the control of the synthetic jets, the near-field vortices continuously shed alternately, while the far-field wake gradually stabilizes.
@@
Moreover, the synthetic jets are positioned near the trailing corner of the square cylinder, resulting in a stabilized wake flow after control in Fig. 14(e). Fig. 14(f) illustrates the placement of the synthetic jets and the streamlines around the square cylinder. Due to the influence of the synthetic jets, a stable and symmetric separation bubble forms on the leeward side of the cylinder, with minimal oscillation observed in the wake region’s streamlines. Under the flow conditions at $Re=500$, the stability of the controlled flow is effectively enhanced when the synthetic jets are positioned near the leading edge. When the synthetic jets are placed near the trailing edge, the controlled flow achieves a stable state without vortex shedding. In the work of Yan et al.Yan et al. (2023), the DRL control framework they designed focuses on reducing $C_{D}$ and $C_{L}$, and the control strategy produces obvious control effects at multiple Reynolds numbers. In our research, we not only focus on reducing the lift and drag coefficients, but also on controlling the stability of the wake region and the detachment of the vortex behind the square cylinder. The control strategy we obtained completely suppresses the shedding of vortices in the square cylinder wake area under the flow of $Re=500$. The first reason for this discrepancy is that, in our research, the agent is capable of observing not only the flow field information around the square cylinder but also, more importantly, the flow field information in the most pulsating regions of the wake field. Li and ZhangLi and Zhang (2022) employed a similar probe design approach and developed a control strategy capable of completely suppressing the Kármán vortex street phenomenon in the wake field of a cylinder. More importantly, near the trailing corners, the synthetic jets, by executing blowing or suction actions, introduce small-scale vortex structures, leading to the formation of stable and symmetric separation vortices behind the square cylinder. At $Re=500$, positioning the synthetic jets near the trailing corners can directly influence the development of the boundary layer, effectively disrupting flow separation and enhancing flow stability.
@@
To comprehensively illustrate the control process of synthetic jets positioned near the leading and trailing corner points on the wake flow field around a square cylinder, we present the case of $Re=500$. We display the contours of instantaneous vorticity from the initiation to the end of the control process. Initially, the activation of the synthetic jets affects only the flow near the square cylinder. We focus on this region initially, and as the control progresses, the influence extends to the wake field, ultimately showcasing the instantaneous vorticity values across the entire computational domain. When the synthetic jets are positioned near the leading edge of the square cylinder, the evolution of the instantaneous vorticity field during the control process is described using five representative time snapshots, as depicted in Fig. 15. In Fig. 15(a), the control action initiates near the leading edge of the square cylinder, causing a slight disruption in the originally continuous vorticity field. Moving to Fig. 15(b), as the synthetic jets near the leading edge continue their control action, their influence on the vorticity around the square cylinder gradually extends trailingward. By the time Fig. 15(c) is reached, vortices generated from flow separation near the leading edge reattach to the cylinder’s sidewalls, elongating along the walls towards the trailing corner of the cylinder. As the flow control progresses to Fig. 15(d), the originally continuous large vortex around the cylinder is disrupted and divided into several smaller vortices. These smaller vortices develop towards the trailing corner, and after separating again, form alternating vortex shedding in the wake region. When the synthetic jets are located near the leading edge of the square cylinder, after the flow control reaches a stable state in Fig. 15(e), alternating vortex shedding occurs on the leeward side of the cylinder, and the shed vortices gradually dissipate in the far field due to the dissipation effect.
@@
III.3.2 Width of Synthetic Jets
Synthetic jets parameters, particularly the width of the jets, significantly influence flow control performance alongside its positioning. In this section, we focus on studying how the width of synthetic jets impacts flow control effectiveness. To achieve the complete suppression of vortices in the wake flow, we position the synthetic jets near the trailing corner of the square cylinder. The width of the synthetic jets under study ranges from $D/20$, $D/25$, $D/30$ to $D/35$, where D represents the characteristic dimension of the square cylinder. Performance metrics for flow control include the reduction rate of the average drag coefficient, and the standard deviations of lift and drag coefficients, as well as the average and standard deviation of action. The stability of training results is evaluated using the standard deviations of $C_{D}$, $C_{L}$, and action. We conduct experiments under two flow conditions, $Re=100$ and $Re=500$, to test the training results for the four jets width. The summarized performance metrics are presented in Tab. 4, and the time-history curves of $C_{D}$, $C_{L}$, and action are illustrated in Fig. 17.
@@ III.3.2 Width of Synthetic Jets
Figure 17: Effect of synthetic jets width on flow control, for $Re=100$ and $Re=500$. $W_{1}$, $W_{2}$, $W_{3}$, and $W_{4}$ represent jets width of $D/20$, $D/25$, $D/30$, and $D/35$, respectively. (a) $Re=100$, $C_{D}$; (b) $Re=500$, $C_{D}$; (c) $Re=100$, $C_{L}$; (d) $Re=500$, $C_{L}$; (e) $Re=100$, Action; (f) $Re=500$, Action.
@@
Tab. 4 presents the results of training the jets at four different widths. At $Re=100$, the average drag coefficient reduction ranges from 13.1% to 15.2% as the width of the synthetic jets transitions from $D/20$, $D/25$, $D/30$ to $D/35$, with the standard deviations of the $C_{D}$ all approaching 99.9%. The impact on the drag reduction rate and the stability of the $C_{D}$ is minor across the four widths of the synthetic jets. However, the reduction in the standard deviation of the $C_{L}$ under $D/30$ and $D/35$ widths is smaller than that under $D/20$ and $D/25$. Further reduction in the width of the synthetic jets adversely affects the stability of the $C_{L}$. The average and standard deviation of actions are greater under $D/30$ and $D/35$ compared to $D/20$ and $D/25$. Decreasing the width of the synthetic jets leads to an increase in both the average and standard deviation of actions utilized. This shows that as the jet width decreases, the control system requires larger amplitude and varying range of actions to achieve the same control effect. At $Re=500$, the reduction rate in the average drag coefficient ranges from 48.9% to 53.0%, with the width of the synthetic jets having minimal impact on the drag reduction rate. However, the reduction rates in the standard deviations of the $C_{D}$ and $C_{L}$ decrease with decreasing width of the synthetic jets. Additionally, the average value and standard deviation of action show consistency with $Re=100$, where narrowing the width of the synthetic jets results in an increase in both the average value and standard deviation of action.
@@
The time curves of $C_{D}$, $C_{L}$, and control actions for the square cylinder at $Re=100$ and $Re=500$ are depicted in Fig. 17. For the $Re=100$ case, at four different synthetic jets widths, the initial stages show significant fluctuations in $C_{D}$, $C_{L}$, and control actions. After this period of intense fluctuation, $C_{L}$ and control actions stabilize near zero, while $C_{D}$ stabilizes near its minimum value. Upon observing the stabilized results, it is evident that the width of the synthetic jets has a minimal impact on drag reduction, although narrower jets width adversely affect the stability of $C_{L}$ and control actions. In the case of $Re=500$, $C_{D}$, $C_{L}$, and control actions also experience significant fluctuations during the initial stages across all four synthetic jets widths. However, the stability of $C_{D}$, $C_{L}$, and control actions is generally worse compared to the $Re=100$ case. This is attributed to the inherently more unstable flow field at $Re=500$, making control more challenging. Comparing the control results for different synthetic jets widths, further narrowing the jet width has a more significant negative impact on the stability of $C_{D}$, $C_{L}$, and control actions.